OpenStudy (anonymous):

Compare the rates of change of the following items. A. The rate of change of item II is greater than the rate of change of item I. B. The rate of change of item I is greater than the rate of change of item II. C. The rate of change of item I is equal to the rate of change of item II.

OpenStudy (anonymous):

https://app147.studyisland.com/pics/218789pic47.png

Nnesha (nnesha):

so what is slope of the equation ??

OpenStudy (anonymous):



Nnesha (nnesha):

that is formula and yes u have to use that to find slope of the graph

Nnesha (nnesha):

y = mx + b slope intercept form equation where m is slope always with x and b is y-intercept

OpenStudy (godlovesme):

@TreyYoung did u find the slopes yet?

OpenStudy (anonymous):

No I Dont Understand How To :(

OpenStudy (godlovesme):

well @Nnesha willl help u :)

Nnesha (nnesha):

for first use that info i gave u above ^^^^^^^

Nnesha (nnesha):

what is m in this equation \[ y = \frac{ -7 }{ 5 }x + 1\]

Nnesha (nnesha):

y = mx + b slope intercept form equation where m is slope always with x and b is y-intercept